AGw Commissioning with LUCIFER 14-15 October 2008

Observers: J. Hill , with LUCIFER folks (N.Ageorges,W.Seifert, et al) and DIMM folks (J.Schmidt,T.Herbst) and A. Rakich and D. Thompson in Tucson
Software Support: N. Cushing in Tucson (came in to office from home)
Telescope Operators: A. Ceranski

Summary

GCS Auto exposure algorithm is not working properly. It should be reviewed with Dave Thompson.

Lucifer reported an image jump see time 4:55 UT.

23:06:25 - 23:06:35 UT - Image jump seen on GCS guide image.

GCS guide hot spot updated to: x=289, y=252. This must be updated in the SVN repository.

Have confirmed that GCS and IDL are acting differently. IDL is collimating and GCS is driving M1 into Y limit. This needs to be fixed ASAP. Doug and Torsten need to work on this.

Dave get luci image data cubes luci_20081014_125 through luci_20081014_136 for Andrew

Details

2:10 UT - Telescope problems being worked on...

4:00 - 4:42 UT - various issues with GCS acquisition, probably a problem with new scaling algorithms in GCS. Tried unsuccessfully to contact Torsten.

Auto exposure needs to work better. The algorithm needs to be reviewed with Dave Thompson.

4:42:52 UT - Reverted to older version of GCS and GCSGUI (GCS.08101301 and GCSGUI.08101301 for both programs) and acquisition now works.

A few times the acquisition image was seen but the guide star was lost moving to center it and when 1st subimage was taken.

5:13:10 to 5:13:30 UT - Lucifer reported a jump in the image.

Had to call John Hill to discuss how to get primary mirror to stop running into Y limit. Andrew thinks this may be the guide star is more than 4 arc-minutes off axis.

5:55 UT - This preset with WFS caused us to hit the Y limit again.

6:06:25 - 23:06:35 UT - Image jump seen on GCS guide image.

6:06:30 UT - We hit Y limit again. Requesting guide stars that are less than 4 arc-minutes off axis.

6:25 UT to 9:30 UT - Wrestled with many issues and experimenting with idl and gcs wfs.

During this time jet had to be reset. Thus had to restart gotemp, PCS, and MCS. Had trouble getting TCS to run properly. Restarting just the subsystem and getting Active presets to work was problematic. It appeared GCS was not running to the point that guiding could start. Attempted to restart all TCS servers without success. Had to completely restart TCS and all servers to get things to work.

Rescued another baby Gecko roaming the building near the 5th floor Coke machine and set it free by the Cherry St front door.

9:40:52 UT - Guiding failed to acquire star when it is obviously was visible in the bottom of the image.

9:57 - Doug/Andrew should look at Image wfsc000028.fits. This is the last image to look at before the hotspot moving experiment.

HOTSPOT EXPERIMENT

10:00 UT - Send in some focus 100 microns with global Z

Run in Guide mode with Doug's idl providing WFS images.

First WFS file is wsfc000027.fits

get hot spot is: -x 289 -y 250 (Original Hot Spot Position)

WFS image is wsfc000028.fits

set hot spot to: -x 289 -y 245

WFS image is wsfc000029.fits

set hot spot to: -x 289 -y 235

WFS image is wsfc000030.fits (Clipping)

Moving back to center (-y 255)

set hot spot to: -x 289 -y 255

WFS image is wsfc000031.fits

set hot spot to: -x 289 -y 265

WFS image is wsfc000032.fits (Maybe a little clipping starting to show).

set hot spot to: -x 289 -y 275

WFS image is wsfc000033.fits (To much Clipping).

set hot spot to: -x 289 -y 270

WFS image is wsfc000034.fits (Clipping).

The offset seems to be clipping on each side at 270 and 235 center is at 252.5 (were at 255 before we started).

Going back to 235 for second look in small steps so as not to lose guiding.

WFS image is wsfc000035.fits

In the M1 PSF GUI changing global offset from 1 Z to .9 Z to reduce the defocus to allow for higher edge cut off resolution.

WFS image is wsfc000036.fits (no vignetting)

Now moving from -y 235 to -y 232 (starting to see

WFS image is wsfc000037.fits (Starting to see some vignetting)

Now moving from -y 235 to -y 230

WFS image is wsfc000038.fits (Starting to see some vignetting)

In the M1 PSF GUI changing global offset from .9 Z to .8 Z to reduce the defocus to allow for higher edge cut off resolution. Essentially focussed.

WFS image is wsfc000039.fits (In focus, no vignetting)

Now moving from -y 230 to -y 228

WFS image is wsfc000040.fits (Hint of vignetting showing on the opposite side going from Z .9 to Z .8)

Now moving from -y 228 to -y 226

WFS image is wsfc000041.fits (Significant vignetting)

Now moving from -y 226 to -y 224

WFS image is wsfc000042.fits (Very significant vignetting)

Now moving from -y 224 to -y 222

WFS image is wsfc000043.fits (More Very significant vignetting, some parts spilling out)

Now moving from -y 222 to -y 220

WFS image is wsfc000044.fits (Some More Very significant vignetting, some parts spilling out)

Now moving from -y 220 to -y 218

WFS image is wsfc000045.fits (Bright Cells = 11, More Some More Very significant vignetting, some parts spilling out)

Now moving from -y 218 to -y 216

WFS image is wsfc000046.fits (Bright Cells = 11, More Some More Very significant vignetting, some parts spilling out)

Now moving from -y 216 to -y 214

WFS image is wsfc000047.fits (Good Enough, Bright Cells = 11, More Some More Very significant vignetting, some parts spilling out)

Updated collimation model.

Taking image wsfc000048.fits, (Looks a bit further out, Cell count increased).

Now moving to -y 218

WFS image is wsfc000049.fits (Close to the edge.)

Check other side.

Now moving to -y 280 in steps not to lose guiding.

WFS image is wsfc000051.fits

Now moving from -y 280 to -y 284

WFS image is wsfc000052.fits (Starting to see some vignetting)

Now moving from -y 284 to -y 288

WFS image is wsfc000053.fits (More vignetting)

Now moving from -y 288 to -y 290

WFS image is wsfc000054.fits (Some More vignetting)

Collimating

WFS image is wsfc000055.fits (Image spots are gone)

Now moving to -y 286

WFS image is wsfc000056.fits (comparible image)

y = (286 + 218) / 2 = 252

Setting y to 252, Moving x

x = 289 y = 252

Moving x to: -x 259 -y 252, in small steps (20 pixels)

Collimate

WFS image is wsfc000057.fits

Now moving from -x 259 to -x 255

WFS image is wsfc000058.fits (Starting to show some vignetting)

Now moving from -x 255 to -x 253

WFS image is wsfc000059.fits (Almost gone)

Moving to other side.

Now moving x to: -x 319

Collimate

WFS image is wsfc000060.fits (Starting to lose light)

Now moving from -x 319 to -x 323

WFS image is wsfc000061.fits (Showing a little more lost cells ?)

Now moving from -x 323 to -x 325

WFS image is wsfc000062.fits ()

x = (325+253) / 2 = 289

HOTSPOT x=289, y=252

11:00 UT - Move back to x=289, y=252 (It took 1 hour to perform the above serious of steps!)

Set zernikes to decollimate telescope to see if GCS can bring it into collimation:

Z7 = +4000 Z4 = +1000

Edit /lbt/tcs/current/etc/GCS/AIP_L.cfg and change guide cam hotspot to x=289, y = 252.

11:18:37 UT - Preset issue in ACTIVE mode

First GCS WFS image since hot spot readjust is wsfc000807.fits

GCS sent big correction for focus and comma. Eventually caused M1 to hit Y limit.

Rolled in Z7 of +1000 to move away from Y limit

Using IDL to collimate.

WFS image is wsfc000064.fits

Sending Zs to primary.

Last image in this series is wsfc000065.fits

Have confirmed that GCS and IDL are acting differently. IDL is collimating and GCS is driving M1 into Y limit. Torsten and Doug need to work together to review and adjust GCS algorithm.

Field Aberration Measurements

11:30 UT - Observing star BS9107

Preset in GUIDE mode and use IDL to perform WFS for collimation with on-axis guide star as a starting point.

Series of through focus images.

First image... WFS image is wsfc000067.fits

Hit Y limit again. Starting over. Moving both M1 and M2 in Y to compensate. M1 = 2.6, M2 = -2.766.

Global offsets entered for Y of M2 = -2.766, M1 = 2.6 and M1 RX = 35.2

Collimate again.

WFS image is wsfc000068.fits sent to Primary. Moved in +Y direction.

What are the signs on M1 and M2 for Y to move them as a rigid body in the Y direction? Chris?

WFS image is wsfc000068.fits sent to Primary. Converging.

To test the sign of Y for M1 and M2, taking M2= 3.766, M1=1.6: Prior wfsc000071 produced Z7 = 123.35

wfsc000072 produced Z7 = -117.43, Andrew upon reflection thinks we were stuck on the Y limit. Would like to discuss with John H. Now thinks signs are the right way.

Collimating again, images wsfc000072.fits - wsfc0000xx.fits

11:58 UT - Preset in guide mode.

Z4 = 0, Lucifer image is luci_20081014_0125 data cube of 10 images, 2 seconds each.

Going to absolute Z4 = +2000, Lucifer image is luci_20081014_0126 data cube of 10 images, 2 seconds each.

Going to absolute Z4 = +1000, Lucifer image is luci_20081014_0127 data cube of 10 images, 2 seconds each.

Going to absolute Z4 = -1000, Lucifer image is luci_20081014_0128 data cube of 10 images, 2 seconds each.

Going to absolute Z4 = -2000, Lucifer image is luci_20081014_0129 data cube of 10 images, 2 seconds each.

Note: Nancy reports seeing stars move in field from data cube to data cube following the application of the new Z4.

Going to absolute Z4 = 0, Lucifer image is luci_20081014_0130 data cube of 10 images, 2 seconds each.

Going to off axis guide star, collimate off axis.

12:09:07 UT - Preset guide mode preset, BS9107, G* = 13, PosAng = ?

Collimated

Z4 = 0, Lucifer image is luci_20081014_0131 data cube of 10 images, 2 seconds each.

Going to absolute Z4 = +2000, Lucifer image is luci_20081014_0132 data cube of 10 images, 2 seconds each.

Going to absolute Z4 = +1000, Lucifer image is luci_20081014_0133 data cube of 10 images, 2 seconds each.

Going to absolute Z4 = -1000, Lucifer image is luci_20081014_0134 data cube of 10 images, 2 seconds each.

Going to absolute Z4 = -2000, Lucifer image is luci_20081014_0135 data cube of 10 images, 2 seconds each.

Going to absolute Z4 = 0, Lucifer image is luci_20081014_0136 data cube of 10 images, 2 seconds each.

12:24:42 UT - Preset guide mode preset, BS9107 (wrong because BS9105 was actually used), G* = 12, PosAng = 306, did not find the star. Resetting pointing, IE=-15.73 CA=-5.445

12:27:49 UT - Preset guide mode preset, BS9107 (wrong because BS9105 was actually used), G* = 12, PosAng = 306, did not find the star.

12:32:44 UT - Preset guide mode, BS9107, WFS on axis, G* = 0. Found the star. Hit the limit.

Global offset, M2 = -2.766, M1 = +2.6, refound the guide star.

Collimating again, wsfc000084.fits - wsfc000089.fits (Converging very slowly).

Global offset, M2 = -3.766, M1 = +1.6, refound the guide star.

Last image had about the same abberation.

Determined Luci team and us have been using different G* lists.

12:46:07 UT - Preset guide mode, BS9107, G* = 12. Guiding and WFS falling apart because of bright sky.

-- JohnHill - 14 Oct 2008
Topic revision: r6 - 19 Oct 2008, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback